How much money should be invested in an account that earns 9% interest, compounded monthly, if you want to have $500,000 after 20 years?
Answer by mananth(16946)

Principal P = x
Amount= 500000
years=n 20
compounded 12 times a year
Rate = 9.00
Amount = P*((n+r)/n)^n

500000 = x *( 1 + r/n )^ n


500000 =P(1+0.090/12)^ 12*20

500000 = P( 1.00750 )^ 240

500000 /( 1.01 )^ 240 = P

P= 83206.42
Investment = $ 83,206.42